### v1

# Diffusion Meets DAgger: Supercharging Eye-in-hand Imitation Learning

Abstract
A common failure mode for policies trained with imitation is compounding execution errors at test time. When the learned policy encounters states that were not present in the expert demonstrations, the policy fails, leading to degenerate behavior. The Dataset Aggregation, or DAgger approach to this problem simply collects more data to cover these failure states. However, in practice, this is often prohibitively expensive. In this work, we propose Diffusion Meets DAgger (DMD), a method to reap the benefits of DAgger without the cost for eye-in-hand imitation learning problems. Instead of collecting new samples to cover out-of-distribution states, DMD uses recent advances in diffusion models to create these samples. This leads to robust performance from few demonstrations. In experiments conducted for non-prehensile pushing on a Franka Research 3, we show that DMD can achieve a success rate of 80% with as few as 8 expert demonstrations, where behavior cloning reaches only 20%. DMD also outperform competing NeRF-based augmentation schemes by 50%.
I Introduction
Imitation learning is an effective way to train robots. However, even when testing policies in environments similar to those used for training, imitation learning suffers from the well-known Compounding Execution Errors problem: small errors made by the learned policy lead the robot to out-of-distribution states causing the learner to make even bigger errors (compounding errors) [56].
One solution to this problem is via manual collection of expert-labeled data on states visited by the learner, a strategy popularly known as Dataset Aggregation or DAgger [56]. In theory, this solution works. However, putting it into practice is challenging from a practical stand point: driving a robot that doesn’t obey your commands is difficult. Many alternatives have been proposed, e.g. [35, 36], but they all require collecting more expert data. In this paper, we pursue an alternate paradigm: automatically generating observations and action labels for out-of-distribution states. By replacing data collection with data creation, we improve the sample efficiency of imitation learning. In fact, data creation was Pomerleau’s original solution to this problem [51]. We revisit, improve, and automate his solution from 30 years ago using modern data-driven image generation methods. We do this for imitation learning problems in context of eye-in-hand setups (i.e. setups where images come from a camera mounted on the robot hand) that are increasingly becoming more popular [64, 84, 77].
Figure 1: Eye-in-hand Imitation learning with DMD: A common failure mode in an imitation learning setting is the problem of poor generalization due to compounding execution errors at test time as shown in (a). This can be solved by collecting more expert data to cover these off-trajectory states as shown in (b) however, this is an expensive process. Our proposed approach is to synthesize data instead of collecting it (c). Magenta arrow represents small perturbation ($\Delta p$) to the trajectory. Cyan arrow represents label ($\tilde{a}_{t}$) for this out-of-distribution observation. We use a state-of-the-art diffusion model to take images $I_{t}$ from expert demonstrations (d, left) and generate realistic off-trajectory images $\tilde{I}_{t}$ (d, right). Note the distance between the grabber and the apple denoted by the green line. This synthetic data augments expert demonstrations for policy learning, leading to more robust policies.
In more detail: given a behavior cloning trajectory, $\tau$ collected from an expert operating the system, we design generative models to synthesize off-distribution states and corresponding action labels. For example, for the object pushing application in Figure 1, this would correspond to generating off-center views of the object from the good expert trajectory $\tau$. Specifically, we learn a function $f(I_{t},\Delta p)$ that synthesizes the observation $\tilde{I}_{t}$ at a small perturbation $\Delta p$ to the trajectory at time step $t$. Knowledge of this $\Delta p$ lets us compute the ground truth action label $\tilde{a}_{t}$ for this out-of-distribution observation. We augment the expert demonstration $\tau$ with multiple off-distribution samples $(\tilde{I}_{t},\tilde{a}_{t})$ along the trajectory. Access to out-of-distribution states and action labels for how to recover from them improves the performance of the task policy.
In this paper, we develop and evaluate this idea in context of eye-in-hand manipulation policies in the real world. While the idea is intuitive, a number of nuances have to be dealt with carefully to get it to work. As we have access to the entire trajectory, the natural choice is to adopt NeRFs [48, 85] to realize the function $f$. While NeRFs work very well for view synthesis in computer vision, we found that they don’t work well for the synthesis task at hand because of the inevitable deformations in the scene as the gripper manipulates the objects. We thus switched to using diffusion models for this synthesis task. This led to good image synthesis even in situations where the scene deformed over the course of manipulation. Computing action labels for these samples presented yet another challenge (Figure 6). The labels should align with making progress towards the goal. We thus investigated different schemes for sampling $\Delta p$’s and computing action labels $\tilde{a}_{t}$.
This paper presents experiments that evaluate the aforementioned design choices to develop a data creation framework to supercharge eye-in-hand imitation learning. Experiments on a physical platform demonstrate the effectiveness of our proposed framework over baseline data augmentation techniques. On a non-prehensile pushing on a Franka Research 3, we show that DMD can achieve a success rate of 80% with as few as 8 expert demonstrations, where behavior cloning reaches only 20%. We also outperform a recently proposed augmentation scheme based on NeRFs [85] by 50%.
II Related Work
II-A Imitation Learning
Behavior cloning (BC), or training models to mimic expert behavior, has been a popular strategy to train robots over the last many decades [50, 51, 58] and has seen a recent resurgence in interest [28, 86]. Ross et al. [56] theoretically and empirically demonstrate the compounding execution error problem (small errors in the learner’s predictions steer the agent into out-of-distribution states causing the learner to make even bigger errors) and propose a Dataset Aggregation (DAgger) strategy to mitigate it. Over the years, researchers have sought to improve the basic BC and DAgger recipe in several ways [25, 4]. [64, 77, 17, 84, 62] devise ways to ease and scale-up expert data collection, while [83] design ways to collect demonstrations in virtual reality. [35, 36] devise ways to simplify DAgger data collection. [14, 61] model the multiple modes in expert demonstrations, while [49] employ a non-parameteric approach. Researchers have also shown the effectiveness of pre-trained representations for imitation learning [49, 74]. [85, 12, 46, 81] employ image and view synthesis models to synthetically augment data. Complementary to our work, [12, 46, 81] focus on generalization across objects via semantic data augmentation by adding, deleting and editing objects in the scene.
[85, 31, 16, 32] pursue synthetic data augmentation to address the compounding execution error problem. [31, 16] conduct these augmentation in low dimensional state spaces. Ke et al. [32] leverage local continuity in the environment dynamics to generate corrective labels for low-dimensional state spaces and LiDAR observations in simulation. Closest to our work, Zhou et al. [85]’s SPARTN model train NeRF’s on demonstration data to synthesize out-of-distribution views. NeRF assumes a static scene. Thus, SPARTN can’t reliably synthesize images when the scene undergoes deformation upon interaction of the robot with the environment. Our use of diffusion models to synthesize images gets around this issue and experimental comparison to SPARTN demonstrate the effectiveness of our design choice.
II-B Image and View Synthesis Models
Recent years have witnessed large progress in image generation [18, 24, 54, 63, 65]. This has led to a number of applications: text-conditioned image generation [54, 5, 20, 57, 52, 82], image and video inpainting [44, 9, 47], image to image translation [27] and generation of novel views for objects and scenes from one or a few images [72, 69, 80, 42, 34]. While diffusion models have proven effective at image synthesis tasks, Neural Radiance Fields (NeRFs) [48] excel at view interpolation tasks. Given multiple (20 – 100) images of a static scene, vanilla NeRFs can effectively interpolate among the views to generate photo-realistic renders. Researchers have pursued NeRF extensions that enable view synthesis from few images [79, 19, 7, 75] and even model deforming scenes [66, 8]. In general, NeRFs are more effective at interpolation problems, while diffusion models excel at problems involving extrapolation (e.g. synthesizing content not seen at all) or where exactly modeling changes in 3D is hard (e.g. deforming scenes). As our work involves speculating how the scene would look if the gripper were at a different location, we adopt a diffusion based methods. We show comparison against a NeRF based method as well.
II-C Diffusion Models and Neural Fields for Robot Learning
Effective generative models and neural field based representations have been used in robotics in other ways than for data augmentation. Researchers have used diffusion models for representing policies [37, 1, 38, 13, 71], generating goals or subgoals [30, 6, 10], offline reinforcement learning [3, 71, 43], planning [40, 29, 73, 41], behaviorally diverse policy generation [23], and predicting affordances [76]. Additionally, they have been used to generate samples with high reward [26] and skill aquisition from task or play data [11]. Neural fields have been used to represent scenes for manipulation [39, 70] and navigation [2].
III Approach
Figure 2: DMD System Overview: Our system operates in three stages. a) A diffusion model is trained, using task and play data, that can synthesize novel views of a scene relative to a given image. b) This diffusion model is used to generate an augmenting dataset that contains off-trajectory views ($\tilde{I_{2}^{1}}$, $\tilde{I_{2}^{2}}$) from expert demonstrations. Labels for these views ( cyan arrows) are constructed such that off-trajectory views will still converge towards task success (right). Images with a green border are from trajectories in the original task dataset. Purple-outlined images are diffusion-generated augmenting samples. c) The original task data and augmenting dataset are combined for policy learning.
Figure 3: DMD Architecture: We use the architecture introduced in [80], a U-Net diffusion model with blocks composed of convolution, self-attention, and cross attention layers. The conditioning image $I_{a}$, and noised target image $I_{b}$ are processed in parallel except at cross-attention layers. The pose conditioning information is injected at cross-attention layers.
III-A Overview
Given task data $\mathbb{D}$, imitation learning learns a task policy $\pi$. The task data takes the form of a collection of trajectories $\tau_{i}$, where each trajectory in turn is a collection of image action pairs, $(I_{t},a_{t})$. $\pi$ is trained using supervised learning to regress action $a_{t}$ from images $I_{t}$ over the task data $\mathbb{D}$. For the non-prehensile manipulation of objects using an eye-in-hand cameras considered in this paper, images $I_{t}$ correspond to views from the eye-in-hand camera as input, and the actions $a_{t}$ are the relative end-effector poses between consecutive time steps.
When trained with few demonstrations, $\pi$ exhibits poor generalization performance. To address poor generalization, our approach generates an augmented dataset $\mathbb{\tilde{D}}$ and trains the policy jointly on $\mathbb{\tilde{D}}\cup\mathbb{D}$. Augmented samples are specifically generated to be out-of-distribution from $\mathbb{D}$ and thus aid the policy’s generalization capabilities. $\mathbb{\tilde{D}}$ is generated through a conditional diffusion model $f$ that generates a $\Delta p$ perturbed view $\tilde{I}$ of an image $I_{t}$ from a given trajectory $\tau$: $\tilde{I}=f(I_{t},\Delta p)$. We use the action labels in the trajectory $\tau$ to compute the action label $\tilde{a}$ for this perturbed view. We describe the design of the conditional diffusion model in Section III-B and detail the augmenting image sampling and action labels computation procedure in Section III-C.
III-B Diffusion Model for Synthesizing $\Delta p$ Perturbed Views
The function $f(I,\Delta p)$ is realized using a conditional diffusion model that is pretrained on large Internet-scale data. Specifically, we adopt the recent work from Yu et al. [80] and finetune it on data from $\mathbb{D}$. The model produces image $I_{b}$ by conditioning on a reference image $I_{a}$ taken by camera $a$ and a transformation matrix ${}_{a}T_{b}$ that maps points in camera $b$’s frame to camera $a$’s frame. By representing the desired $\Delta p$ as the transformation between two cameras (${}_{a}T_{b}=\Delta p)$, and using images from $\mathbf{D}$ as the reference images ($I_{a}\sim\mathbf{D}$), the learned diffusion model can generate the desired perturbed views for augmenting policy learning.
III-B1 Model Architecture
A diffusion model can be thought of as an iterative denoiser: given a noisy image $x_{t}$, diffusion timestep $t$, it is trained to predict the noise $\epsilon$ added to the image.
$\mathcal{L}=\|\epsilon-\epsilon_{\theta}(x_{t},t)\|^{2}_{2}$
It is typically realized via a U-Net [55]. Our application requires conditional image generation where the conditions are source image $I_{a}$ and transformation ${}_{a}T_{b}$. Thus, by using $I_{b}$ as the diffusion target, and conditioning on $I_{a}$ and ${}_{a}T_{b}$, the model $\epsilon_{\theta}$ learns to denoise new views of the scene in $I_{a}$ based on the transformation ${}_{a}T_{b}$. Finally, rather than directly denoising in the high-resolution pixel space, the denoising is done in the latent space of a VQ-GAN autoencoder, $E$ [15, 53]. This gives the final training objective of:
$\mathcal{L}=||\epsilon-\epsilon_{\theta}(x^{b}_{t},E(I_{a}),{}_{a}T_{b},t)||\quad\text{where}\quad x^{b}_{0}=E(I_{b}).$
Following [80], the pose conditioning information ${}_{a}T_{b}$ is injected into the U-Net via cross-attention as shown in Figure 3.
III-B2 Model Training
Figure 4: Visualization of perturbed images generated by diffusion model for augmenting policy learning. In each row, we show $\tilde{I_{t}}$ generated from $I_{t}$ through different camera translations.
Training the model requires access to triples: $(I_{a},I_{b},{}_{a}T_{b})$. We process data from $\mathbb{D}$ to generate these triples. Let each trajectory consist of images $I_{1},\dots,I_{N}$. We use the structure from motion [21, 68] implementation from the COLMAP library [59, 60] to extract poses for the images in the trajectory $\tau$. This associates each image $I_{t}$ in the trajectory with the (arbitrary) world frame ${}_{t}T_{w}$. We can then compute relative pose between arbitrary images $I_{a}$ and $I_{b}$ from a trajectory via ${}_{a}T_{b}={}_{a}T_{w}{}_{w}T_{b}$. We sample random pair of images from the trajectory to produce triples $(I_{a},I_{b},{}_{a}T_{b})$ that are used to train the model $\epsilon_{\theta}$. COLMAP also outputs the camera intrinsics necessary for training the model. Since COLMAP reconstruction is only correct up to a scale factor, we normalize translations by the largest displacement between adjacent frames per trajectory.
As $\mathbb{D}$ contains much less data than is typically required to train a diffusion models from scratch, we finetune the model from Yu et al. [80]. We find that finetuning with as few as 24 trajectories leads to realistic novel view synthesis for our task as shown in Figure 4. We only finetune the diffusion model weights and use the pre-trained VQ-GAN codebooks. We experiment with different datasets for finetuning this diffusion model: task data from $\mathbb{D}$, play data [45, 78] collected in our setup, and their combinations.
III-C Generating Out-of-Distribution Images and Labels
III-C1 Sampling Out-of-distribution Images
We generate out of distribution views for images in the task dataset $\mathbf{D}$. $\Delta p$’s are sampled as follows. For an image $I_{t}$ from trajectory $\tau$, we sample vectors in a random direction, with magnitude drawn uniformly from $[0.2s,s]$, where $s$ is the largest displacement between adjacent frames in $\tau$. $\Delta p$ simply corresponds to translation along this randomly chosen vector. We constrain these samples to not be too aligned with the actual action.
III-C2 Labeling Generated Images
Figure 5: Training Examples from the Diffusion Model: We visualize generated examples, $\tilde{I}$, used to train our policies along with the computed action label (the arrow is a projection of the 3D action into the 2D image plane: the arrow pointing up means move the gripper forward, pointing to the right means move it right). We show augmentations generated by our diffusion model for at different stages of the task. The first row shows augmentations from images where the grabber approaches the apple, while the second row shows augmentations from images where the grabber moves the apple towards the goal location.
Figure 6: The overshooting problem: Under certain conditions, the inferred action for $\tilde{I}_{t}$ may direct the agent away from task success. We refer to this as the overshooting problem. At time step $t+1$, the view $I_{t+1}$ in (c) has moved to the right from $I_{t}$ in (a). However, the synthesized sample $\tilde{I}$ (b) has moved even further to the right than $I_{t+1}$. In (b), Cyan arrow represents action label for $\tilde{I}_{t}$ computed using $I_{t+1}$ as the target; green arrow represents action label computed using $I_{t+3}$ as the target. Since $\tilde{I}_{t}$ has “overshot” $I_{t+1}$, an action taken with $I_{t+1}$ as the next intended target moves backward, away from the goal. Computing the action with respect to a farther image, say $I_{t+3}$ in (d), does not have this issue.
For each image $\tilde{I}_{t}=f(I_{t},\Delta p)$ generated from an original image $I_{t}$, we use $I_{t+k}$ as the target for generating the action label. The action label for $\tilde{I}_{t}$ is simply the action that would convey the agent from the pose depicted in $\tilde{I}_{t}$ to the pose in $I_{t+k}$.
From COLMAP, we obtain camera pose information ${}_{t}T_{w}$ for time step $t$ and ${}_{t+k}T_{w}$ for time step $t+k$. The synthesized view $\tilde{I}_{t}$ can be treated as an image taken by a virtual camera whose pose is represented as ${}_{t}T_{\tilde{t}}$. The diffusion model synthesizes perturbed view conditioned on $I_{t}$ and ${}_{t}T_{\tilde{t}}\;$, or $\tilde{I}_{t}=f(I_{t},{}_{t}T_{\tilde{t}})$ (Section III-B). The action label $\tilde{a}_{t}$ for $\tilde{I}_{t}$ is then computed as $\tilde{a}_{t}={}_{\tilde{t}}T_{t+k}\;=\;{}_{\tilde{t}}T_{t}\,{}_{t}T_{w}\,({}_{t+k}T_{w})^{-1}$. Examples of $(\tilde{I}_{t},\tilde{a}_{t})$ are shown in Figure 5.
One might wonder why we don’t just use $k=1$? As the scale of each trajectory is different, the diffusion model learns the distribution of scales over $\mathbb{D}$. Depending on the scale of the COLMAP reconstruction, this causes the generated samples to move more than 1 unit for some trajectories. When using $k=1$, some generated samples move past the target image $I_{t+1}$ (as described in Figure 6). This causes conflicting supervision, as the computed action does not make progress toward completing the task. Using a large $k$ guards against this, leading to stronger performance downstream, as our experiments verify (Section IV-C2).
Figure 7: Robotic Platform for Online Testing: We conduct our experiments on a Franka Research 3 robot with a wrist-mounted GoPro Hero 9. In the depicted configuration, the task is to push the apple onto the red circle. Since VIME’s [77] action space is relative 3D translations, we modified its grabber mount on the robot, allowing the robot to move without reaching joint limits while preserving the original rotation of the end-effector.
IV Experiments
IV-A Experiment Setup
IV-A1 Task, Observation Space, Action Space
We use the non-prehensile pushing task from VIME [77] as a benchmark task. The task is to push the target object to a target location marked by a red circle and requires both reaching and pushing the object. Following Young et al. [77], we use a grabber as the robot end-effector as shown in Figure 7. This allows for easy collection of expert demonstrations using a GoPro mounted on a Grabber. Due to the difference in joint limits between the XArm (used by VIME) and the Franka Research 3 robot, we redesign the attachment between the grabber and the robot flange. In this way, the robot can execute planar push in the typical top-down configuration, extending the workspace to a wider range of the table. The comparison is shown in Figure 7. Observations come from the eye-in-hand GoPro camera. We use the action space used in VIME’s [77] publicly available code: relative 3D translations of the end-effector.
IV-A2 Task and Play Data
We collect task and play data trajectories using a GoPro mounted on a Grabber. For task data, an expert pushes the object to a target location. For play data, we move the grabber around the apple and push it in many different directions. Through this procedure, we only get image sequences. We use structure from motion [21, 68] to extract camera poses ${}_{t}T_{w}$ for images in the collected sequences. Action $a_{t}$ is computed as the relative translation between $I_{t+1}$ and $I_{t}$ in $I_{t}$’s camera frame. Since COLMAP only gives reconstructions upto an unknown scale factor, following [77] we normalize the action to unit length and interpret it as just a direction. Since the final robot action space is 3D translation, we avoid rotating the grabber during both type of data collection; however, slight rotation of the grabber still exists in expert demonstrations.
IV-A3 Policy Architecture and Execution
We adopt the architecture for the task policy $\pi$ from VIME [77] but replace the AlexNet [33] backbone with an ImageNet pre-trained ResNet-18 backbone [22]. The policy consists of the ResNet backbone followed by 3 MLPs that predict the action. This task policy accepts a center cropped image from the GoPro camera and outputs the direction in which the camera should move. The policy is trained by L1 regression to the (unit length) actions using the Adam optimizer with a learning rate of 1e-4. Actions are executed on the robot by commanding the robot to go 1cm in the predicted direction.
IV-A4 Comparisons
We use behavior cloning on the expert data as done in past work [77], as the baseline. We refer to this as BC. We also compare to SPARTN [85] and evaluate the various design choices. Since there is no publicly available code from SPARTN, we replicate their procedure using NerfStudio [67].
IV-B Visual Comparison of Generated Images
We start by assessing the visual quality of samples generated by the diffusion model in DMD. Figure 4 and Figure 5 shows sample images generated by our diffusion model. The model is able to faithfully generate images where the camera is moved left / right and front / back with minimal artifacts.
We also compare the visual quality of our generated images to those generated using the NeRF based approach from SPARTN [85]. The moving gripper in the scene breaks the static scene assumption made by NeRF. Thus, [85] mask out the gripper before training the NeRF. While this works for the pre-grasp trajectory that Zhou et al. [85] seek to imitate, this strategy fails when the gripper actually manipulates the scene, as in our task. Thus, we investigate different masking schemes: a) no masking, b) masking a larger region around the gripper, c) masking just the gripper; and show visualizations in Figure 8. No masking leads to the worst results as expected. Masking just the grabber fixes the grabber, but causes the object being manipulated to blur out. Masking a larger region around the grabber fixes the blurring but also eliminates the object being manipulated. In contrast, images synthesized using our diffusion model move the camera as desired without creating any artifacts.
Figure 8: Diffusion vs NeRF Here we visualize perturbed samples generated using DMD and NeRF baselines with different masking strategies. The top row shows images generated corresponding to a forward movement relative to $I_{t}$, and the bottom row for a backward movement. When using a NeRF with no masking (a), the reconstructions of the gripper and apple are degenerate, as they violate the static scene assumption in NeRFs. With a box-mask (b), the apple is occluded by the mask, and consequently, missing from the reconstruction. Even with a mask around the end-effector (Grabber-Mask), there are major artifacts in reconstructing the apple (c). With ours (Diffusion Model), the model faithfully reconstructs the grabber and apple (d).
IV-C Offline Validation
We collected 74 expert demonstrations in total for the apple pushing task. We use 24, 23, and 27 trajectories for train, validation, and test, respectively. We use vanilla behavior cloning on the expert data as the baseline as done in past work [77]. Our offline validation involves evaluating the policy’s predictions on the test set. While the test set samples doesn’t contain examples of going off the distribution, it does measure generalization to some extent by evaluating the policy on states outside of the training set. For the error metric, we use the median angle between the predicted and the ground truth translation.
We conduct offline validation to compare with other data augmentation schemes and select the value of $k$ to mitigate the over-shooting problem. Online experiments in Section IV-D test the success rate of the policy.
Table I: Comparisons between diffusion-generated augmented dataset and standard augmentation schemes. The table shows median angle error (in radians) between predicted and ground-truth translations. Adding $\mathbb{\tilde{D}}$ generated by diffusion models improve performance on top of other augmentation techniques.
Methods BC DMD
No Aug 0.376 0.349
w/ Color jitter 0.381 0.350
w/ Flip 0.356 0.338
w/ Color jitter & Flip 0.355 0.328
IV-C1 Comparison with other data augmentation schemes
We want to understand the effectiveness of augmenting with diffusion-generated images compared to augmenting with standard techniques, such as color jitter and horizontal flip. Table I shows that augmentation, in general, improves performance, and augmenting with out-of-distribution images leads to larger improvement compared to standard techniques. Comparing the two columns, we see that adding diffusion-generated images improve performance in all settings.
IV-C2 Overshooting problem, what frame to use?
Figure 9: Effect of using different future frames for labeling augmenting images on translation prediction: to overcome the overshooting problem described in Figure 6, we experiment with different future frame $I_{t+k}$ for labeling the diffusion model generated images. Error decreases as $k$ increases and plateaus around $k=3$. Therefore, we use frame $I_{t+3}$ for labeling $\tilde{I_{t}}$
In order to mitigate the overshooting problem, we use image $I_{t+k},k>1$ to label the augmenting image instead of $I_{t+1}$. We evaluate policies trained with labels produced from different $k$’s, and show the results in Figure 9. We find that labeling with $I_{t+1}$’s pose works worse than labeling with pose for $I_{t+3}$ and beyond. As the curve plateaus beyond $k=3$ and using larger $k$ leads to the exclusion of the last $k-1$ frames in each trajectory, we use $k=3$ for our online experiments.
Table II: Online Robot Experiment Results. We conduct 5 experiments: 10 trials per method and randomize the apple start locations between trials.
IV-D Online Validation
Figure 10: Comparison of different methods on staying on course. We show the trajectories executed by BC, SPARTN [85], and DMD over several steps towards the target. Unlike BC and SPARTN, which gradually deviate from the intended path, DMD maintain its course towards the right and reach the target successfully. See project website for more execution videos.
Our online experiments measure the success rate at the non-prehensile pushing task. We conduct 10 trials per method and randomize the apple start location between trials. To prevent human biases, we choose the apple start location before sampling the method to run next. Execution trajectories can be found on the project website.
IV-D1 DMD vs. Behavior Cloning
We take the best model for BC and DMD from Table I and evaluate them on the robot. We find the vanilla BC model to get a 30% success rate while DMD achieves 100% (Table IV-C2 a). This significant difference shows that minor error at each step can accumulate and result in task failure.
IV-D2 DMD with Few Demonstrations
Motivated by the 100% performance of DMD with 24 demonstrations, we ask how low can the number of demonstrations be? We consider 2 low data settings with 8 and 16 demonstrations. For these settings, the diffusion model is trained with only play data. Success rate for these two models on the robot is shown in Table IV-C2 c. Once again DMD outperforms BC and achieves 80% success even when trained with only 8 demonstrations.
IV-D3 DMD vs. SPARTN [85]
Following up on the visual comparison of the quality of the augmenting samples from Section IV-B, we conduct a head-to-head comparison against SPARTN on the real robot. SPARTN achieves a success rate of 50% vs. DMD succeeds 100% of the time (Table IV-C2 b). In Figure 10, we show that while SPARTN diverges from the expert course and pushes the apple off the table, DMD brings the apple to the target sucessfully.
IV-D4 Utility of Play Data for Training Diffusion Model
We also experimentally validate what the choice of the dataset necessary for training an effective diffusion model for view synthesis. We repeat the experiment with 24 demos but vary the diffusion model that generated the augmenting samples. Using the diffusion model trained on just task data lead to an 80% success rate vs. 100% success rate with the diffusion model trained on a combination of task and play data (Table IV-C2 d). Thus, while just training the diffusion model on task data is effective (a nice result in practice), performance can be boosted further by training the diffusion model on play data.
IV-D5 Scaling to Other Objects
We conduct an experiment on another object, a cup. The BC baseline achieves a success rate of 80% vs DMD achieves a success rate of 90% (Table IV-C2 e).
V Discussion
DMD was designed to improve the sample efficiency of eye-in-hand imitation learning. Compared to traditional behavior cloning, DMD shows significant improvement in experiments involving non-prehensile pushing across 2 objects. Remarkably, DMD provides effective augmenting strategies using generated images and enabled learned policies to achieve 80% success rate from as few as 8 demonstrations. DMD not only performs better than conventional augmentation schemes used in robotics but also shows improved metrics when combined with them.
Our experiments revealed the benefits of the different parts of DMD over past approaches. The use of diffusion models to synthesize images allows DMD to generate augmenting images for manipulation tasks that involve non-static scene. Prior work [85] used NeRFs instead of diffusion models, restricting generation to samples from pre-grasp part of the trajectories. We showcase the advantages of diffusion models over NeRFs through qualitative results in Figure 8 and real robot experiments in Section IV-D3. We also tested which type of data: task data, play data, or a combination of both, should be used for training the diffusion model. We found all versions to outperform behavior cloning. This suggests two practically useful results. First, DMD can directly be used to improve a task policy without the need to collect any additional data, i.e. the task data itself can be used to finetune a novel-view synthesis diffusion model to augment the task dataset to improve performance. Second we can benefit from play data, which maybe simpler to collect than task data. To foster replication and future work, we will make code, data, and models publicly available.
Acknowledgments
We give special thanks to Kevin Zhang for 3D printing the attachment for us. This material is based upon work supported by the USDA/NSF AIFARMS National AI Institute (USDA #2020-67021-32799) and NSF (IIS-2007035).